A True-False test has 10 questions with each having 2 possible answers with one correct. Assume a student randomly guesses the answers to every question.
A. What is the probability of getting exactly 5 correct answers?
B. What is the probability of getting less than 6 correct answers?
This is what I have come up with so far n=10 p=2 =.05 for p
p(5) + P(10) =.000+
P(6) + P(10) =.000+

Answer by stanbon(57347)

A True-False test has 10 questions with each having 2 possible answers with one correct. Assume a student randomly guesses the answers to every question.
----------
P(correct) = 1/2 ; P(wrong) = 1/2
------------------------------------
A. What is the probability of getting exactly 5 correct answers?
Binomial Problem with n = 10, p = 1/2 , x = 5
P(x=5) = 10C5*(1/2)^5*(1/2)^5 = 252*0.00097656 = 0.2461
------------------------------------------------------------------
B. What is the probability of getting less than 6 correct answers?
P(0<= x <=5) = binomcdf(10,1/2,5) = 0.6230..
----------------------------------------------------------
Cheers,
Stan H.